INTRACAVITY LASER SPECTROSCOPY OF NICKEL CHLORIDE: SYSTEM H

\underline{JAMES J. O’BRIEN}, Department of Chemistry, University of Missouri, St. Louis, MO 63121-4499; LEAH O’BRIEN, Department of Chemistry, Southern Illinois University, Edwardsville, IL 62026-1652.

The near infrared electronic transition of NiCl occurring in the region of 12,259 cm^{−1} has been recorded by intracavity laser absorption spectroscopy. The NiCl molecules were produced in a nickel hollow cathode operating with an applied potential of 350 to 750 V, using 0.5 to 2.2 torr argon or helium, and a trace amount of carbon tetrachloride. At this time we believe it does connect to the X ^{2}Π_{3/2} state. A high resolution analysis of this transition is in progress and results of the fit will be presented.